# CycleGAN-Driven Transfer Learning for Electronics Response Emulation in High-Purity Germanium Detectors

### v1

@@ 1 Introduction
High-Purity Germanium (HPGe) detectors play a crucial role in numerous low-background experimental searches for neutrinoless double-beta decay and dark matter [1, 2, 3, 4, 5]. In these experiments, the HPGe detectors used often have a point contact geometry, which features a small readout electrode, called a point contact, at one end of the crystal insulated from the outer surface which is held at high voltage. This layout lowers the detector capacitance, leads to relatively long charge drift times (of $\mathcal{O}(\mu s)$) and yields a very steep gradient in the weighting-field near the point contact. In combination, these features produce pulse shapes that are highly sensitive to the energy deposition position, allowing for multi-site and surface event discrimination based on detector pulse shapes. These techniques are broadly termed pulse-shape discrimination (PSD). A variant of the point contact design is the inverted-coaxial point contact (ICPC) detector, which combines the low capacitance and excellent energy resolution of point-contact detectors with a larger active volume typical of coaxial designs. A cutaway view of an ICPC detector is shown in the left panel of figure 1.
The microphysics in the bulk of HPGe detectors is very well understood, allowing for accurate simulation of pulses. Charges, freed into the valence band by an energy deposition, drift under the electric field to the readout electrodes, inducing the measured signal. Pulse shape simulation (PSS) pipelines using publicly available tools such as fieldgen and siggen [6] or SolidStateDetectors [7] software packages start from an electric field simulation that takes into account the geometry and impurity concentration of each detector. Each energy deposition taken from a particle interaction simulation tool such as Geant4 [8, 9, 10] is used to simulate the pulse induced by charge drift in the electric field and collection from the deposition site. The pulses are scaled according to the energies deposited at each site and summed together to create the total event pulse.
@@
PSS are critical to designing high-performing detectors, understanding event topology, and distinguishing between signal and background events. [11, 12]. PSS aims to generate pulses based on these parameters that are indistinguishable from detector pulses as measured in data. This allows us to produce ground-truth datasets where a bijective mapping is established between each simulated pulse and its corresponding incident particle(s). Suppose we apply a pulse shape discrimination cut that is commonly used to identify and reject background events in $0\nu\beta\beta$ experiments to this dataset, (see Ref. [13]). In that case, it will be immediately apparent which event topologies lead to the signal sacrifice or background contamination of this cut. This information can be used for evaluation and further improvement of cut efficiency. Pulse shape simulation is also critical for designing new cuts and new HPGe detector geometries, as well as training machine learning models to replace or supplement traditional discrimination methods.
@@
Current experiments avoid detailed electronics modeling by calculating reconstruction features directly from Monte Carlo particle-interaction simulations, relying on heuristic methods to meet most of their simulation needs [14]. These methods can be tuned to match experimental data in a particular region of interest, but generally do not reproduce pulse shape behavior correctly for all spectral features. They also fall short in their ability to account for detector-by-detector variations and operational changes in experimental conditions, which can evolve over time.
Transfer learning can help overcome the need for a first-principles electronics model and parameter retuning by directly learning the translations between each simulated pulse and its corresponding detector pulse. In this work, we present a novel transfer learning framework called Cyclic Positional U-Net (CPU-Net) for improving pulse shape simulation. CPU-Net translates simulated pulses into detector data-like pulses so that they closely resemble real detector pulses. We demonstrate that CPU-Net improves the reproduction of key pulse shape features, such as current amplitude, rise times, and tail slopes, without including these characteristics as targets of the network training.
@@ 2.1 Mathematical Formalism
The traditional pulse shape simulation approach typically learns $P(X^{\prime}|Y^{\prime})$ and applies it to a given $Y$ in $\mathcal{D}_{Source}$ so that the set of simulated pulses $(\mathcal{X})$ matches the set of data pulses $\mathcal{X}^{\prime}$. For example, one can fit the tail slope of the detector pulses to determine the decay constant and add a pole with that time decay to the simulation to replicate the data. This method requires a complicated collection of modeling and characterization data, along with computationally expensive fitting procedures in a highly degenerate high-dimensional parameter space [15, 16].
@@
We intentionally do not impose an explicit loss on the target-domain feature distribution $P(Y^{\prime})$. Constraining the network solely at the waveform level ensures it remains agnostic to any particular set of pulse parameters. This allows newly devised pulse-shape based analysis techniques to be studied without bias. Moreover, the electronic response is a global effect that should correct features common to all pulses. By not fitting $P(Y^{\prime})$ directly, the translator avoids compensating for event-by-event physics differences, allowing residual disagreements between simulations and data to serve as clean diagnostics of the systematic uncertainty associated with PSS methods. In short, the goal of CPU-Net is to perform a translation of each simulated waveform rather than generating entirely new pulses.
2.2 Positional U-Net
@@
Our method relies on a U-Net [17], a convolutional neural network initially developed for biomedical image segmentation, as a baseline model when designing the REN. A U-Net contains contracting (encoding) and expansive (decoding) channels. The contracting path contains $n$ convolutional layers to encode a pulse into a characteristic vector to capture contextual information and reduce the spatial resolution of the input. The expanding path contains $n$ upsample layers to decode the feature vector back to an output with the same length. The network structure also allows information to flow at different levels to the decoding part to provide maximum reconstruction efficiency. The U-Net structure is depicted in figure 2, where the tensor shape is also denoted at each stage. The Conv1d module in the U-Net is a series of layers; within each layer, the kernel size is an important hyperparameter to control the receptive field of convolutional layers, and padding is added to guarantee the same input and output shapes. Max pooling is used in all Conv1d modules except for the first layer to reduce the feature map size. This increases the effective receptive field by focusing on the most essential features in the signal.
Figure 2: Left: Layer-wise breakdown of the Positional U-Net. The blue lines represent the skip connections between contracting and expanding paths of the U-Net. Positional encoding layers are also shown for all levels. Right: Cycle-consistent adversarial training with PU-Net as the Response Emulation Network (REN) and Inverse Response Emulation Network (IREN). The red oval region represents simulated pulses in $\mathcal{T}_{Source}$, and the blue oval region represents detector pulses in $\mathcal{T}_{Target}$. REN transfers simulated pulses to data-like pulses, while the IREN transfers data pulses to simulation-like pulses. A RNN discriminator differentiates data-like pulses outputs from the REN from true data. Similarly, another RNN discriminator differentiates simulation-like pulses outputs from IREN from true simulations.
@@
During network design, we observed that the conventional U-Net does not reproduce the tail of the pulse due to the lack of positional information in intermediate layer outputs. Therefore, we developed a Positional U-Net (PU-Net) model with layer-wise positional encoding maps $\mathcal{M}_{\mathrm{position}}$ inspired by the Transformer model [18]. $\mathcal{M}_{\mathrm{position}}$ contains sine and cosine functions with different frequencies. Since each U-Net layer outputs a tensor with a different shape, $\mathcal{M}_{\mathrm{position}}$ must be generated separately for each layer and added to the layer output, as shown in the left panel of figure 2.
@@
Finally, we added a reparameterization trick to the bottom of PU-Net. The trick was developed by Kingma and Welling [19] to sample a random space while preserving the gradient flow. Reference [20] pointed out that this trick has the ability to increase the stochasticity of machine learning models. Based on our experiments, increased stochasticity helps the model learn the distributions of the reconstruction features better. We call this model Positional U-Net (PU-Net). Together, the PU-Net model has $7,213,781$ trainable parameters.
2.3 RNN with Attention
@@
A RNN is a natural discriminator for pulses, because it can intrinsically preserve the temporal ordering of a waveform. We chose an attention-coupled recurrent neural network (RNN)[21] as the discriminator for CPU-Net. A single-layer bidirectional RNN is used as the discriminator model in CPU-Net. We adopt the Gated Recurrent Unit (GRU) [22] as the internal structure of the RNN. The raw input pulses are first embedded in the $m=128$ space. Embedding is done by initializing a lookup table that maps integer ADC values to dense vectors of fixed size. These vector representations are learned during training iterations to capture meaningful pulse features. We use an embedding trick that optimizes computation by directly retrieving the corresponding row from the embedding matrix instead of performing unnecessary matrix multiplications.
@@
The embedded pulse is fed into a GRU that processes it sequentially while maintaining hidden states to capture temporal dependencies. This yields a $64$-dimensional output $\vec{I}(t)$ at each intermediate step $t$ as well as a final output $\vec{F}$. We then use an attention mechanism [21] which can boost the performance of RNN by allowing it to focus on different parts of the pulses, such as the rising edge during different steps of the training. The attention mechanism contains an attention matrix $A$ of dimension ($64$,$64$), which is used to calculate the attention score between $\vec{F}$ and each $\vec{I}(t)$. This calculates how much each time step contributes to the final output:
@@ 2.4 CycleGAN
Ideally, the REN should be trained with paired pulses of simulations and data. In reality, collecting such a paired dataset ranges from a major challenge to functionally impossible, depending on the desired population of events. Although we can easily obtain $Y$ from $X$ through simulation, obtaining $Y^{\prime}$ from $X^{\prime}$ is difficult without precise knowledge of $P(X^{\prime}|Y^{\prime})$. This means it is impossible to obtain a paired dataset between $\mathcal{D}_{Source}$ and $\mathcal{D}_{Target}$.
The Cycle Generative Adversarial Network (CycleGAN) framework [23, 24] provides an unsupervised learning approach to train the REN using adversarial losses. We first construct two networks: a REN  $\Lambda$ and an Inverse REN $\bar{\Lambda}$, both with PU-Net structure. We then construct two discriminator networks $\delta_{S}$ and $\delta_{T}$ for the source and target pulses, respectively, each based on a RNN with attention. We call this model, which combines all the features, the Cyclic Positional U-Net (CPU-Net). The trained CPU-Net produces both a REN and an IREN, translating pulses between the source and target domains. The REN is the primary interest of this work, but the IREN could be used to improve data analysis methods.
@@ 3.1 Data selection
The LEGEND Collaboration characterized subsets of newly produced HPGe detectors at Oak Ridge National Laboratory (ORNL). For this study, we selected characterization data from a LEGEND Inverted-Coaxial Point-Contact (ICPC) detector, V06643A, manufactured by ORTEC [25]. In the ORNL setup, the detector was mounted in a vendor cryostat, shielded by lead, and housed in a concrete alcove to reduce the external background. A Thorium-$228$ source was placed on top of the cryostat. The resulting detector signals were digitized using a FlashCam digitizer, and the pygama package [26] and dspeed package [27] were used to convert the FlashCam output to LEGEND HDF5 files. The event energies were calibrated to convert the ADC units to keV. The energies were then corrected for charge trapping. Figure 3 shows the final energy spectrum obtained from the detector.
@@
Given that the FEP has high statistics and contains a mixture of single- and multi-site events, the FEP provides an excellent sample of events for training. SEP and DEP can then be used as validation data sets, evaluating the model’s response to multi-site and single-site events, respectively. DEP and SEP are not perfectly pure samples since bremsstrahlung and Compton events introduce a continuum of background events. However, they are dominated by single-site and multi-site topologies, respectively, offering appropriate populations to validate the performance of CPU-Net.
Figure 3: Left: Calibrated energy spectrum from a 228Th source at ORNL compared to the Monte Carlo simulation results. Key peaks from 228Th source are labeled. The simulation does not include detector surface effects, nor any electronic-noise or energy resolution effects, thus resulting in a mismatch. Right: The simulation geometry of the ORNL characterization setup, showing the detector in green, the source in red, the aluminum cryostat and holder in dark grey, and the lead shielding in light grey.
@@ 3.2 Simulations
Geant4 was used to geometrically simulate events in the ORNL characterization setup. The geometry modeled in Geant4 is shown in figure 3. It includes a HPGe detector, a radioactive source, and an Aluminum PopTop cryostat holder surrounded by lead shielding. We simulated $100$ million 228Th decay events originating from the source and recorded their energy depositions within the HPGe detector.
@@
Given the location of energy deposition within the detector, we used siggen simulations to generate pulses for hits in the DEP, SEP, and FEP energies. The siggen simulations treat every energy deposit as a single point charge and heuristically account for diffusion and self-repulsion by convolving the output signal with a Gaussian that mimics the effects of a charge cloud with 0.1 mm full-width at half-maximum (FWHM). We disabled the preamplifier decay in the simulation. The simulation does not include detector surface effects, which can degrade event energies in regions within $1$-$2$ mm of the detector contact surfaces. Electronic noise and other detector-level effects affecting the energy resolution are also not included in the simulations. These cause spread in the measured spectrum due to statistical fluctuations in charge carrier statistics, recombination, and electronic noise. The simulated energy spectrum was obtained by energy-weighted summation of the pulses for energy depositions for a given event. Figure 3 also shows the simulated spectrum.
3.3 Post processing and Network Training

### v2

@@ 1 Introduction
High-Purity Germanium (HPGe) detectors play a crucial role in numerous low-background experimental searches for neutrinoless double-beta decay and dark matter [LEGEND_2025, MJD, GERDA, SuperCDMS, COGENT]. In these experiments, the HPGe detectors used often have a point contact geometry, which features a small readout electrode, called a point contact, at one end of the crystal insulated from the outer surface which is held at high voltage. This layout lowers the detector capacitance, leads to relatively long charge drift times (of $\mathcal{O}(\mu s)$) and yields a very steep gradient in the weighting-field near the point contact. In combination, these features produce pulse shapes that are highly sensitive to the energy deposition position, allowing for multi-site and surface event discrimination based on detector pulse shapes. These techniques are broadly termed pulse-shape discrimination (PSD). A variant of the point contact design is the inverted-coaxial point contact (ICPC) detector, which combines the low capacitance and excellent energy resolution of point-contact detectors with a larger active volume typical of coaxial designs. A cutaway view of an ICPC detector is shown in the left panel of figure 1.
The microphysics in the bulk of HPGe detectors is very well understood, allowing for accurate simulation of pulses. Charges, freed into the valence band by an energy deposition, drift under the electric field to the readout electrodes, inducing the measured signal. Pulse shape simulation (PSS) pipelines using publicly available tools such as fieldgen and siggen [icpc_siggen] or SolidStateDetectors [ssd_paper] software packages start from an electric field simulation that takes into account the geometry and impurity concentration of each detector. Each energy deposition taken from a particle interaction simulation tool such as Geant4 [geant4one, geant4two, geant4three] is used to simulate the pulse induced by charge drift in the electric field and collection from the deposition site. The pulses are scaled according to the energies deposited at each site and summed together to create the total event pulse.
@@
PSS is critical to designing high-performing detectors, understanding event topology, and distinguishing between signal and background events. [MERTENS201981, Comellato_2020ljj]. PSS aims to generate pulses based on these parameters that are indistinguishable from detector pulses as measured in data. This allows us to produce ground-truth datasets where a mapping is established between each simulated pulse and its corresponding incident particle(s). Suppose we apply a pulse shape discrimination cut that is commonly used to identify and reject background events in $0\nu\beta\beta$ experiments to this dataset, (see Ref. [AvsE]). In that case, it will be immediately apparent which event topologies lead to the signal sacrifice or background contamination of this cut. This information can be used for evaluation and further improvement of cut efficiency. Pulse shape simulation is also critical for designing new cuts and new HPGe detector geometries, as well as training machine learning models to replace or supplement traditional discrimination methods.
@@
Current experiments avoid detailed electronics modeling by calculating reconstruction features directly from Monte Carlo particle-interaction simulations, relying on heuristic methods to meet most of their simulation needs [legend_pcdr]. These methods can be tuned to match experimental data in a particular region of interest, but generally do not reproduce pulse shape behavior correctly for all spectral features. They also fall short in their ability to account for detector-by-detector variations and operational changes in experimental conditions, which can evolve over time.
Transfer learning can help overcome the need for a first-principles electronics model and parameter retuning by directly learning the translations between each simulated pulse and its corresponding detector pulse. In this work, we present a novel transfer learning framework called Cyclic Positional U-Net (CPU-Net) for improving pulse shape simulation. CPU-Net translates simulated pulses into detector data-like pulses so that they closely resemble real detector pulses. We demonstrate that CPU-Net improves the reproduction of distributions of key pulse shape parameters, such as current amplitude, rise times, and tail slopes, without including these distributions or the parameter values as targets of the network training.
@@ 2.1 Mathematical Formalism
The traditional pulse shape simulation approach attempts to approximate $P(X^{\prime}|Y^{\prime})$ by introducing nuisance parameters into $P(X|Y)$ and fitting $\mathcal{T}_{Target}$ to obtain their values. These parameters are applied to a given $Y$ in $\mathcal{D}_{Source}$ so that the set of simulated pulses $(\mathcal{X})$ matches the set of data pulses $\mathcal{X}^{\prime}$. For example, one can fit the tail slope of the detector pulses to determine the decay constant and add a pole with that time decay to the simulation to replicate the data. This method requires a complicated collection of modeling and characterization data, along with computationally expensive fitting procedures in a highly degenerate high-dimensional parameter space [Ben_Thesis, Sam_Thesis].
@@
We intentionally do not impose an explicit loss on the target‐domain feature distribution $P(Y^{\prime})$. Constraining the network solely at the waveform level ensures it remains agnostic to any particular set of pulse parameters. This allows newly devised pulse-shape based analysis techniques to be studied without bias. Moreover, the electronic response is a global effect that should correct features common to all pulses. By not fitting $P(Y^{\prime})$ directly, the translator avoids compensating for event‑by‑event physics differences, allowing residual disagreements between simulations and data to serve as clean diagnostics of the systematic uncertainty associated with PSS methods. In short, the goal of CPU‑Net is to perform a translation of each simulated waveform rather than generating entirely new pulses.
A limitation of this approach is that electronics chain introduces stochastic effects and other effects that destroy information, e.g. electronic noise and limited bandwidth at high frequency. Therefore, an invertible mapping between simulation and data does not exist. Regardless of this limitation, this method provides a very good approximation for practical application, as the pulse shape parameters of interest are designed to be as insensitive to noise and high-frequency effects as possible.
2.2 Network Architecture
2.2.1 Positional U-Net
@@
Our method relies on a U-Net [UNet], a convolutional neural network initially developed for biomedical image segmentation, as a baseline model when designing the REN. A U-Net contains contracting (encoding) and expansive (decoding) channels. The contracting path contains $n$ convolutional layers to encode a pulse into a characteristic vector to capture contextual information and reduce the spatial resolution of the input. The expanding path contains $n$ upsample layers to decode the feature vector back to an output with the same length. The network structure also allows information to flow at different levels to the decoding part to provide maximum reconstruction efficiency. The U-Net structure is depicted in figure 2, where the tensor shape is also denoted at each stage. The Conv1d module in the U-Net is a series of layers; within each layer, the kernel size is an important hyperparameter to control the receptive field of convolutional layers, and padding is added to guarantee the same input and output shapes. Max pooling is used in all Conv1d modules except for the first layer to reduce the feature map size. This increases the effective receptive field by focusing on the most essential features in the signal.
Figure 2: Left: Layer-wise breakdown of the Positional U-Net. The blue lines represent the skip connections between contracting and expanding paths of the U-Net. Positional encoding layers are also shown for all levels. Right: Flow of pulses in the RNN+Attention discriminator. The embedded waveforms pass through a bidirectional GRU and attention layer to form context-aware features. The output is concatenated with hidden states from the GRU. This is fed to a fully connected network which produces a classification decision using a sigmoid function.
Figure 3: A flowchart representing the cycle-consistent adversarial training of the Cycle-GAN network. The solid lines indicate the flow of training instances through the PU-Net Response Emulation Network (REN) and Inverse Response Emulation Network (IREN) generators, and dotted lines represent how each stage is used to produce a subset of the losses used to train the network. The full list of the losses used is given in Table 1. The red lines represent the forward-cycle, which begins with simulated pulses and transfers them to data-like pulses using the REN. The Data Discriminator $\delta_{D}$ differentiates data-like pulses from true data and is used for adversarial training of the REN. Completing the cycle through the IREN to produce recovered simulation pulses allows the use of an a weighted mean absolute error (L1) loss to ensure cycle consistency. The reverse cycle, shown by the blue lines, is trained simultaneously with the same strategy, beginning with data pulses.
@@
During network design, we observed that the conventional U-Net does not reproduce the tail of the pulse due to the lack of positional information in intermediate layer outputs. Therefore, we developed a Positional U-Net (PU-Net) model with layer-wise positional encoding maps $\mathcal{M}_{\mathrm{position}}$ inspired by the Transformer model [vaswani2017attention]. $\mathcal{M}_{\mathrm{position}}$ contains sine and cosine functions with different frequencies. Since each U-Net layer outputs a tensor with a different shape, $\mathcal{M}_{\mathrm{position}}$ must be generated separately for each layer and added to the layer output, as shown in the left panel of figure 2.
@@
Finally, we added a reparameterization trick to the bottom of PU-Net. The trick was developed by Kingma and Welling [VAE] to sample a random space while preserving the gradient flow. Reference [AAE] pointed out that this trick has the ability to increase the stochasticity of machine learning models. Based on our experiments, increased stochasticity helps the model learn the distributions of the reconstruction features better. We call this model Positional U-Net (PU-Net). Together, the PU-Net model has $7,213,781$ trainable parameters.
2.2.2 RNN with Attention
@@
A RNN is a natural discriminator for pulses, because it can intrinsically preserve the temporal ordering of a waveform. We chose an attention-coupled recurrent neural network (RNN)[bahdanau2015neural] as the discriminator for CPU-Net; its overall structure is shown in the right panel of figure 2. A single-layer bidirectional RNN is used as the discriminator model in CPU-Net. We adopt the Gated Recurrent Unit (GRU) [cho2014learning] as the internal structure of the RNN. The raw input pulses are first embedded in the $m=128$ space. Embedding is done by initializing a lookup table that maps integer ADC values to dense vectors of fixed size. These vector representations are learned during training iterations to capture meaningful pulse features. We use an embedding trick that optimizes computation by directly retrieving the corresponding row from the embedding matrix instead of performing unnecessary matrix multiplications.
@@
The embedded pulse is fed into a GRU that processes it sequentially while maintaining hidden states to capture temporal dependencies. This yields a $64$-dimensional output $\vec{I}(t)$ at each intermediate step $t$ as well as a final output $\vec{F}$. We then use an attention mechanism [bahdanau2015neural] which can boost the performance of RNN by allowing it to focus on different parts of the pulses, such as the rising edge, during different steps of the training. The attention mechanism contains an attention matrix $A$ of dimension (6464,$64$), which is used to calculate the attention score between $\vec{F}$ and each $\vec{I}(t)$. This calculates how much each time step contributes to the final output:
@@ 2.3 CycleGAN
Ideally, the REN would be trained with paired pulses of simulations and data. In reality, collecting such a paired dataset ranges from a major challenge to functionally impossible, depending on the desired population of events. Although we can easily obtain $Y$ from $X$ through simulation, obtaining $Y^{\prime}$ from $X^{\prime}$ is difficult without precise knowledge of $P(X^{\prime}|Y^{\prime})$. This means it is impossible to obtain a paired dataset between $\mathcal{D}_{Source}$ and $\mathcal{D}_{Target}$.
The Cycle Generative Adversarial Network (CycleGAN) framework [CycleGAN, hoffman2018cycada] provides an unsupervised learning approach to train the REN using adversarial losses. We first construct two networks: a REN $\Lambda$ and an Inverse REN $\bar{\Lambda}$, both with PU-Net structure. We then construct two discriminator networks $\delta_{S}$ and $\delta_{T}$ for the source and target pulses, respectively, each based on a RNN with attention. We call this model, which combines all the features, the Cyclic Positional U-Net (CPU-Net). The overall structure of the CycleGAN training framework is depicted in figure 3; a detailed description of the framework, losses, and training is given in section 3.4. The trained CPU-Net produces both a REN and an IREN, translating pulses between the source and target domains. The REN is the primary interest of this work; future work will study whether the IREN can be used to improve data analysis methods.
@@ 3.1 Data Selection
The LEGEND Collaboration characterized a subset of newly produced HPGe detectors at Oak Ridge National Laboratory (ORNL). For this study, we selected characterization data from a LEGEND Inverted-Coaxial Point-Contact (ICPC) detector, V06643A, manufactured by ORTEC [ortec]. In the ORNL setup, the detector was mounted in a vendor cryostat, shielded by lead, and housed in a concrete alcove to reduce the external background. A Thorium-$228$ source was placed on top of the cryostat. The resulting detector signals were digitized using a FlashCam digitizer, and the pygama package [pygama] and dspeed package [dspeed] were used to convert the FlashCam output to LEGEND HDF5 files. The event energies were calibrated to convert the ADC units to keV. The energies were then corrected for charge trapping [mjd_pole_zero]. Figure 4 shows the final energy spectrum obtained from the detector.
@@
The training losses are expectations over the empirical training samples and, therefore, depend on the choice of the training set. We intentionally train on the high-statistics Full Energy Peak (FEP), which contains a mixture of single-site and multi-site topologies. This choice enables the network to learn a translation that is common across topologies rather than fitting topology-specific physics. To test this generalization, we validate it on the SEP (multi-site–dominated) and the DEP (single-site–dominated) regions of the spectrum. These are not perfectly pure samples, as Compton scattering introduces an underlying continuum of background events, and bremsstrahlung creates a population of multi-site DEP events. However, they are dominated by single-site (in the DEP) and multi-site (in the SEP) topologies, offering appropriate populations to validate the performance of CPU‑Net.
Figure 4: Left: Calibrated energy spectrum from a 228Th source at ORNL compared to the Monte Carlo simulation results. Key peaks from 228Th source are labeled. The simulation does not include detector surface effects, nor any electronic‑noise or energy resolution effects, thus resulting in a mismatch. Right: The simulation geometry of the ORNL characterization setup, showing the detector in green, the source in red, the aluminum cryostat and holder in dark grey, and the lead shielding in light grey.
@@ 3.2 Simulations
Geant4 was used to simulate particle interactions in the ORNL characterization setup. The geometry modeled in Geant4 is shown in figure 4. It includes a HPGe detector, a radioactive source, and an Aluminum PopTop cryostat holder surrounded by lead shielding. We simulated $100$ million 228Th decay events originating from the source and recorded their energy depositions within the HPGe detector.
@@
Given the Geant4 simulations of energy depositions within the detector, we used siggen simulations to generate pulses for hits in the DEP, SEP, and FEP energies. The siggen simulations treat every energy deposit as a single point charge and heuristically account for diffusion and self-repulsion by convolving the output signal with a Gaussian that mimics the effects of a charge cloud with 0.1 mm full-width at half-maximum (FWHM). We disabled the preamplifier decay in the simulation. The simulation does not include detector surface effects, which can degrade event energies in regions within 11-22 mm of the detector contact surfaces. Electronic noise and other detector-level effects affecting the energy resolution are also not included in the simulations. These cause spread in the measured spectrum due to statistical fluctuations in charge carrier statistics, recombination, and electronic noise. The simulated energy spectrum was obtained by energy-weighted summation of the pulses for energy depositions for a given event. Figure 4 also shows the simulated spectrum.
3.3 Post Processing